OpenStudy (anonymous):

Solve the system using elimination. x + 5y = –28 4x – 5y = 13 (–5, –3) (–2, –5.2) (-3, -5) (5, 4)

OpenStudy (anonymous):

add and get \[5x=-15\] then solve that one

OpenStudy (loser66):

Let me interpret what he said: "add and get " means x + 5y = -28 4x -5y =13 --------------- 5x = -15

OpenStudy (loser66):

Now divide both sides by 5 to get x plug back one of the equation to get y. dat sit

OpenStudy (anonymous):

the first answer im confused

OpenStudy (loser66):

hey, -15/5 = -5??? if you say the first answer is the correct one, hehehe

OpenStudy (anonymous):

ha ha I said I think but was not sure lol

OpenStudy (loser66):

-15/5 =??

OpenStudy (anonymous):

3 lol

OpenStudy (loser66):

-3 , not 3

OpenStudy (loser66):

and how many options have -3 as the first number?

OpenStudy (loser66):

you have only 1 option, that is (-3, -5) , no choice for you. hehehe

OpenStudy (anonymous):

so -3 -5 then lol

OpenStudy (loser66):

I always cheat on that way, hehehe, save time.

OpenStudy (anonymous):

now I am lost lol

OpenStudy (loser66):

why?

OpenStudy (anonymous):

because u said -3 -5 and I thought it was the top answer

OpenStudy (loser66):

(-3, -5) means x =-3, y =-5 while the first answer is (-5,-3) , that is x = -5, y = -3 They are different, right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

The solving process gives you x =-3, hence the first one is not the answer, right?

OpenStudy (anonymous):

correct

OpenStudy (loser66):

so?? only option C gives you (-3,.....) I don't care what ..... is, just x =-3, then pick it

OpenStudy (anonymous):

but I want the right answer lol

OpenStudy (loser66):

That is what I said "I cheat on it", instead of continuing solving for y, I pick it and 100% it is correct.

OpenStudy (anonymous):

ok im going with c then